OpenStudy (anonymous):

Line f contains the points (5, –4) and (4, –6). What is the slope of a line perpendicular to it?

OpenStudy (mandre):

Can you calculate the slope of your line?

OpenStudy (mandre):

Line f that is

OpenStudy (anonymous):

no how?

OpenStudy (mandre):

\[slope = \frac{ y _{1}-y _{0} }{ x _{1}-x _{0} }\]

OpenStudy (anonymous):

a. 2 b. 1/2 c. -1/2 d. –2 those r the choices they give me

OpenStudy (mandre):

Just plug in your coords you got into that slope formula.

OpenStudy (anonymous):

how though

OpenStudy (anonymous):

after getting dis slope...use the formula m1*m2=-1

OpenStudy (mandre):

(5, –4) and (4, –6). (x1, y1) and (x0, y0) Got that?

OpenStudy (anonymous):

yaaa

OpenStudy (anonymous):

where m1 and m2 are the slopes of line and the line perpendicular to it

OpenStudy (mandre):

So what's the slope of f?

OpenStudy (anonymous):

have u calculated the first slope?

OpenStudy (anonymous):

im so confused omfg

OpenStudy (anonymous):

ok u r given two points put it in the formula

OpenStudy (mandre):

(5, –4) and (4, –6). (x1, y1) and (x0, y0) x1 = 5, y1 = -4 x0 = 4, y0 = -6

OpenStudy (anonymous):

(-6-(-4))/(4-5) solve it

OpenStudy (anonymous):

-5.5

OpenStudy (anonymous):

-10

OpenStudy (anonymous):

thenn

OpenStudy (anonymous):

-6+4/-1 =-2/-1 =2

OpenStudy (anonymous):

negative multiped by negative is positive

OpenStudy (anonymous):

ohhhhhhhhhhhhhhh

OpenStudy (anonymous):

right ya thanks omg ur a life savior

OpenStudy (anonymous):

now put this m1 value in, m1*m2=-1 u will get it